Now do an Amortization Schedule

Bank A's offer is 15 years at 3.5% interest compounded monthly.

Determine your monthly payment. Write in a complete sentence. Fill in what you would put on the calculator to show work.

Calculator

APPS / Finance / TVM Solver

N = 180

I% = 3.5

PV = -1175000

PMT = Solve

FV = 0

P/Y = 12

C/Y = 12

PMT: end/begin

Alpha Solve

My monthly Payments would be $8,399.87

What is the total amount that you will pay?

Total amount paid = Monthly payment × Total months

                              = $8,399.87 × 180

                              The total I would pay would be $1,511,976.60

What is the total interest you will pay? (Show work):

$1,511,976.60 - $1,175,000

                          The total interest I will pay would be $336,976.60
Answer #1

Using financial calculator

Set mode to END
N=12*15=180
I/Y=3.5%/12
PV=-1175000
FV=0
CPT PMT=8399.87

Monthly payments will be 8399.87 each for 15 years 0r total 180 payments

Total amount paid=Monthly payment*Total number of payments=8399.87*180=1511976.60

Total amount paid will be 1511976.60

Total interest paid=Total amount paid-Loan=1511976.60-1175000=336976.60

Total interest paid will be 336976.60